OpenStudy (rock_mit182):

Find the rational exponent. • Simplify each expression using the rules of exponents and examine the steps you are taking.

OpenStudy (rock_mit182):

\[(4^{16})^{\frac{ 1 }{ 2 }}\]

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ (a^b)^c=a^{b \times c} }\)

OpenStudy (rock_mit182):

I think, first off product rule between exponents

OpenStudy (solomonzelman):

your a=4, b=16 c=1/2

OpenStudy (rock_mit182):

or i can use the radical property ?

OpenStudy (jdoe0001):

\(\Large (4^{16})^{\frac{ 1 }{ 2 }}\implies 4^{\cancel{ 16 }\cdot \frac{ 1 }{ \cancel{ 2 } }}\)

OpenStudy (rock_mit182):

\[\sqrt{4^{18}}\]

OpenStudy (rock_mit182):

i meant 2 ^18

OpenStudy (jdoe0001):

hmm well.. the 18... ... thought it was 16? =)

OpenStudy (jdoe0001):

\(\bf 16\cdot \frac{ 1 }{ 2 }\implies \frac{ 16 }{ 2 }\)

OpenStudy (rock_mit182):



OpenStudy (jdoe0001):

ohh... yes

OpenStudy (rock_mit182):

is that called radical property ¡

OpenStudy (rock_mit182):

?

OpenStudy (jdoe0001):

hmm I'd think so......

OpenStudy (jdoe0001):

\(\bf \large (4^{16})^{\frac{ 1 }{ 2 }}\implies (2^{2})^{\frac{ 16 }{ 2 }}\implies \sqrt{(2^2)^{16}}\implies \sqrt{(2^{16})^2}\)

OpenStudy (rock_mit182):

Ok i will search out each property in order to write out the paper

OpenStudy (rock_mit182):

That's very helpful thanks a lot

OpenStudy (jdoe0001):

yw

OpenStudy (rock_mit182):

\[a ^{\frac{ m }{ n }} = \sqrt[n]{a ^{m}}\]

OpenStudy (rock_mit182):

\[\sqrt{4^{16}}\]

OpenStudy (rock_mit182):

\[\sqrt{4^{16}}= 4^{\frac{ 16 }{ 2 }}=4^{8}\]

OpenStudy (rock_mit182):

\[4^{8} = (2*2)^{8}\]

OpenStudy (rock_mit182):

\[(2)^{^{8}}(2)^{8} = 2^{8+8}\]

OpenStudy (rock_mit182):

\[2^{16}\]